### v1

# Approximating the universal thermal climate index using sparse regression with orthogonal polynomials

Abstract
This article explores novel data-driven modeling approaches for analyzing and approximating the Universal Thermal Climate Index (UTCI), a physiologically-based metric integrating multiple atmospheric variables to assess thermal comfort. Given the nonlinear, multivariate structure of UTCI, we investigate symbolic and sparse regression techniques as tools for interpretable and efficient function approximation. In particular, we highlight the benefits of using orthogonal polynomial bases-such as Legendre polynomials-in sparse regression frameworks, demonstrating their advantages in stability, convergence, and hierarchical interpretability compared to standard polynomial expansions. We demonstrate that our models achieve significantly lower root-mean squared losses than the widely used sixth-degree polynomial benchmark-while using the same or fewer parameters. By leveraging Legendre polynomial bases, we construct models that efficiently populate a Pareto front of accuracy versus complexity and exhibit stable, hierarchical coefficient structures across varying model capacities. Training on just 20% of the data, our models generalize robustly to the remaining 80%, with consistent performance under bootstrapping. The decomposition effectively approximates the UTCI as a Fourier-like expansion in an orthogonal basis, yielding results near the theoretical optimum in the $L_{2}$ (least squares) sense. We also connect these findings to the broader context of equation discovery in environmental modeling, referencing probabilistic grammar-based methods that enforce domain consistency and compactness in symbolic expressions. Taken together, these results illustrate how combining sparsity, orthogonality, and symbolic structure enables robust, interpretable modeling of complex environmental indices like UTCI - and significantly outperforms the state-of-the-art approximation in both accuracy and efficiency.
Keywords:
Climate Index Sparse regression Orthogonal polynomials
1 Introduction
The increasing availability of high-resolution environmental data—driven by advances in remote sensing, sensor networks, and global monitoring—has created new opportunities and challenges for climate analysis [26, 45, 46, 47, 48, 49, 50]. These complex, spatiotemporal datasets offer insights into atmospheric, hydrological, and ecological processes [28, 35, 41, 51, 53], but often exceed the capacity of traditional methods. Machine learning (ML) provides tools well-suited for extracting structure from such data [22, 59, 64], with applications ranging from forecasting to anomaly detection [33, 34]. One emerging use of ML is in refining climate indices—simplified metrics like the Universal Thermal Climate Index (UTCI), Palmer Drought Severity Index (PDSI), or Oceanic Niño Index (ONI)-that summarize complex climate phenomena [24]. The UTCI is a comprehensive bioclimatic index developed to assess the thermal environment with direct relevance to human thermal physiology. The UTCI is expressed in degrees Celsius and is intended to be interpreted analogously to air temperature, though with a more nuanced and physiologically grounded foundation.
The UTCI uses a detailed thermo-physiological model, see Fig. 1, to simulate human thermal responses—like sweating, shivering, and skin blood flow-under changing conditions. This enables more accurate assessments than traditional indices. UTCI categorizes thermal stress on a continuous scale from “extreme cold” to “extreme heat,” aiding public health communication and risk management. Widely used in biometeorology, urban planning, and occupational safety, it supports studies on urban heat islands, outdoor labor, and heat-health warnings. Its standardized approach allows international comparisons and long-term climate analysis, informing strategies for thermal comfort, public safety, and climate adaptation. It is also increasingly used in many countries, especially in Europe, as a measure of thermal comfort for outdoor conditions and is calculated as part of the operational meteorological forecast [15].
Figure 1: Schematic of the UTCI framework [31]: meteorological inputs and a clothing model feed into a physiological model to compute thermal stress. The resulting UTCI value is classified on a color scale from extreme cold to extreme heat stress.
The dataset [11] we use provides tabulated UTCI values computed using the full thermo-physiological model, covering a wide range of environmental variables: air temperature, wind, incoming radiation, and humidity. Formally, the UTCI is defined as: UTCI = Ta + Offset, where Ta is the air temperature and the Offset is the physiologically equivalent temperature difference, representing how other environmental factors modify the effect of the thermal stress on the human body. To approximate this data, the authors also introduced a sixth-degree polynomial regression model, which has since become a standard in human biometeorology due to its simplicity and computational efficiency. The goal of this study is to develop a more accurate and robust model for the UTCI Offset-one that reduces average and extreme errors, enhances interpretability, and offers fast computation for large-scale or real-time applications. Specifically, we explore the use of sparse and symbolic machine learning to approximate UTCI from meteorological inputs. Using orthogonal polynomial bases (e.g., Legendre polynomials), we develop interpretable, numerically stable models that generalize well and offer structural advantages over standard polynomial regressions.
The UTCI Offset accounts for the effects of incoming radiation (expressed via mean radian temperature Tr), of wind speed (va), which influences convective heat loss, and relative humidity (rH), that affects evaporative heat loss (e.g., sweating efficiency). Furthermore, the Offset depends on Ta indirectly (since the effects of wind and humidity depend on it). In Fig. 2(a) we see how the UTCI Offset varies along the different environmental variables. Instead of the humidity (rH), the water vapor pressure (pa) can be used which is a nonlinear function of rH and the air temperature (Ta). However, the variables have different distribution, see Fig. 2(b), which impacts the extent that approximations of UTCI can generalize, discussed below.
(a) (a)
(b) (b)
Figure 2: (a) 3D plot of UTCI Offset [11] at 5% relative humidity, showing how wind speed (va), air temperature (Ta), and mean radiant temperature difference (Tr–Ta) combine to influence thermal stress. Color indicates the UTCI Offset magnitude across these environmental dimensions. (b) The different distributions of the water vapor pressure and relative humidity in the computed Offset dataset [11]. The water vapor pressure is strongly peaked at zero, while the relative humidity is uniform across its range.
2 Methods
Equation discovery aims to learn interpretable mathematical expressions, either differential or algebraic equations, that describe dynamical systems [61, 67, 66]. Positioned at the intersection of symbolic machine learning and system identification, it is increasingly relevant in environmental and climate science, where data-driven yet transparent models are essential. Traditional models rely on expert-derived formulations [36, 37, 38, 39, 40], but the complexity and volume of climate data call for automated alternatives. Symbolic regression, which iteratively combines mathematical operators and variables to fit data, forms the core of equation discovery [60, 9, 63, 13, 16]. These methods are commonly implemented using evolutionary or other (e.g., enumerative) search algorithms [55, 54, 27, 44, 57, 6, 30].
Recent advances integrate probabilistic grammars to incorporate prior knowledge and constrain the search to physically meaningful expressions [8, 21, 7, 29, 32]. This structured approach improves both model interpretability and search efficiency, especially in domains governed by established scientific principles. Equation discovery has been applied to environmental systems [1, 2, 3, 4, 5], including ecosystem dynamics [17, 18, 19, 20, 23, 14, 42, 43, 56, 58], where it can match or surpass expert models while revealing new relationships [62, 65, 68]. Its ability to generate compact, interpretable models makes it especially suitable for climate applications, where model transparency and physical plausibility are vital.
To simplify UTCI computation, the standard approach in the literature is to use a sixth-degree regression polynomial [11, 25]. However, this method suffers from generalization issues [10]. Fig. 3(a) shows the approximation error at 5% relative humidity, while 3(b) displays a histogram of the errors, revealing a normal distribution centered at zero, indicating minimal bias. We aim to improve upon this standard approximation using equation discovery and sparse regression methods. The standard approximation uses the Offset dataset [11] and we employ the same data in our computations, see Fig. 2(a).
(a) (a)
(b) (b)
Figure 3: (a) The difference between the standard UTCI approximation and the real values for relative humidity of 5%. (b) Histogram of the differences (or errors) showing a normal distribution centered at zero.
Sparse machine learning models aim to construct parsimonious predictive functions by enforcing zero-valued coefficients in high-dimensional parameter spaces, thereby performing implicit feature selection [12]. This sparsity promotes interpretability, reduces overfitting, and improves computational tractability, especially when the number of candidate predictors is large or when strong correlations exist among inputs. Sparse regression, a key instantiation of this paradigm, extends linear regression with an $L_{1}$-norm regularization term—most notably in the Lasso (Least Absolute Shrinkage and Selection Operator)—to penalize unnecessary parameters and induce a compact representation.
In this work, we employ sparse regression to identify compact, interpretable models of the UTCI, emphasizing its suitability for high-dimensional input spaces with redundant or weakly relevant features. While sparse modeling is well-established in statistical learning, its application to orthogonal polynomial bases-particularly in the context of bioclimatic indices—remains unexplored. By leveraging the structure of orthogonal polynomials, we obtain improved numerical stability and additive expansions that facilitate coefficient interpretability. To our knowledge, this is the first application of sparse regression using orthogonal bases to approximate the UTCI, addressing both predictive accuracy and model parsimony. Our results show that this approach surpasses the standard sixth-degree polynomial approximation in both accuracy and efficiency.
3 Results and discussion
Table 1 presents a detailed comparison of model performance across a range of polynomial degrees for both standard (non-sparse) linear regression and sparse regression techniques, evaluated in the context of approximating the UTCI. The standard approximation is a sixth-degree regression polynomial model with four variables, consisting of 210 terms and achieving a root mean squared (RMS) loss of 1.12. This serves as the benchmark to be matched or improved upon. It is important to note that the standard approximation [11] does not directly employ the relative humidity (rH), but the water vapor pressure (pa) which is a highly nonlinear function derived from the relative humidity (rH) and air temperature (Ta). As we noted above, the relative humidity is well represented across its entire range, see Fig. 2(b), while the water vapor pressure is strongly peaked around zero. Optimization employing the water vapor pressure (pa) as an independent variable (instead of rH) is thus poorly conditioned and leads to instability in the regression coefficients, both in simple and sparse regression. While using the pa (instead of rH) can achieve better accuracy (lower loss), it comes at the price of losing parameter consistency across optimizations with different polynomial degrees. For this reason we report our results employing the relative humidity (rH) instead of the water vapor pressure (pa).
Method
Standard
1.12
210
Polynomial degree
4th 6th 8th 10th 12th 14th 16th
Linear
regression
2.1
70
1.3
210
0.92
495
0.67
0.71
1001
0.54
0.62
1820
0.44
0.66
3060
0.36
1.74
4845
Sparse
regression
2.1
65
1.38
124
1.03
176
0.88
210
0.69
355
0.63
400
0.6
424
Table 1: Root mean squared train loss, test loss and the number of parameters in approximating the UTCI Offset. The baseline reference, labeled as “Standard,” corresponds to the sixth-degree regression polynomial model with four variables. Unless otherwise stated the test loss equals the train loss. Where two loss values are reported, they indicate a notable train-test discrepancy, typically suggesting overfitting. Training is done with 20% of the data and testing is performed with 80%. Results are robust under bootstrapping.
The regression methods are applied to polynomial basis expansions of increasing degree, evaluated on the basis of root mean squared (RMS) test loss and number of active parameters. Unlike many studies in the literature where models are trained on the majority of the data and evaluated on a relatively small test set, our approach inverts this paradigm: training is conducted on only 20% of the available data, while performance is assessed on the remaining 80%. Despite this stringent evaluation setting, the models achieve comparable performance on both training and test sets, underscoring their strong generalization capabilities. This performance stability is further validated through bootstrapping, which reveals minimal variance in both loss metrics and selected features across resampled datasets. The reported performance metrics—such as train/test loss and number of parameters—remain stable when the model training and evaluation process is repeated on multiple random re-samplings (bootstrapped subsets) of the data. This suggests that the results are not sensitive to specific data splits and that the models generalize well across different subsets of the dataset, indicating reliability and consistency in the reported findings. These findings demonstrate the robustness and reliability of the proposed framework.
Linear regression without any sparsity constraints shows improved performance at higher degrees, with test loss reducing as model capacity increases. However, this comes with a dramatic increase in the number of parameters; it reaches over 1800 coefficients by degree 12. Furthermore, the discrepancy between train and test losses at higher degrees (e.g., 0.62 vs. 0.54 at degree 12) indicates overfitting, despite the improved predictive accuracy. The resulting models are also substantially more complex, raising concerns regarding interpretation and generalization. Sparse regression with standard polynomial bases show similar performance at low degrees but fails to converge beyond the 6th degree. This indicates that enforcing sparsity in a poorly conditioned basis becomes increasingly difficult as model complexity grows.
Figure 4: Loss versus number of parameters for different polynomials degrees. The regularization parameter was varied in the lasso regression to yield a Pareto front in model accuracy and complexity for each degree.
In contrast, sparse regression using an orthogonal Legendre basis exhibits superior stability and accuracy across all degrees. It outperforms the baseline 6th-degree polynomial fit from degree 8th onward, achieving a test loss of 0.88 at degree 10 with only 210 parameters—the same count as the original benchmark model, but with improved generalization. As the degree increases to 16, the loss reduces further to 0.60 using 424 parameters—a fraction of those used by the corresponding standard regression model. The orthogonality of the Legendre basis likely contributes to better numerical conditioning, facilitating sparse model discovery even at high degrees. These results emphasize the importance of basis selection and regularization strategy in symbolic regression tasks. Sparse methods, when combined with well-structured bases like Legendre polynomials, offer a promising path toward accurate, compact, and interpretable models in high-dimensional settings.
Furthermore, optimization of nonlinear objective functions using gradient-based algorithms can be computationally intensive, especially in high-dimensional spaces where convergence is slow and local minima may hinder performance. In contrast, the regression-based approach proposed in this article—particularly through sparse regression with orthogonal polynomials—offers significantly faster computation. By framing the problem as a structured regression task rather than a nonlinear optimization, the method avoids costly iterative procedures and scales efficiently with dimensionality, making it highly suitable for rapid modeling of complex environmental indices like the UTCI.
(a) (a)
(b) (b)
Figure 5: Parameters (or polynomial coefficients) and how they change for different polynomial degrees for (a) simple regression and (c) sparse regression (using Legendre basis).
Fig. 4 illustrates the relationship between model complexity (measured by the number of parameters) and prediction accuracy (log-scaled loss) for sparse regression models using Legendre polynomial bases of varying degrees. Each curve corresponds to a fixed polynomial degree, ranging from 4 to 16, with points reflecting models of increasing complexity obtained through regularization. A clear trend is observed: for a given polynomial degree, increasing the number of parameters generally results in improved model accuracy (i.e., lower loss). However, diminishing returns set in, and the rate of improvement flattens. More notably, the envelope formed by the lowest loss at each level of complexity across all degrees traces an emergent Pareto front. This front captures the trade-off between model simplicity and predictive performance.
Higher-degree models (e.g., degrees 12–16) dominate this frontier at higher parameter counts, offering better loss with only marginal increases in complexity. In contrast, lower-degree models saturate quickly, highlighting their limited expressivity. The Pareto front thus reflects the optimal set of models that balance accuracy and sparsity, guiding model selection under complexity constraints. The use of Legendre polynomials ensures numerical stability and encourages efficient basis representations, which supports the recovery of compact yet accurate models in this sparse regression setting.
The Fig. 5(a) and (b) we visualize the behavior of regression coefficients obtained from simple regression and sparse regression with orthogonal Legendre polynomials. Both plots use a logarithmic x-axis to indicate the parameter index and reveal how coefficients evolve as higher-degree polynomial terms are introduced. In Fig. 5(a), each line corresponds to simple regression solutions using polynomial bases of increasing degree. The x-axis denotes the index of polynomial terms (sorted or sequential), while the y-axis shows the corresponding coefficient values. A key observation is that the coefficients of lower-degree terms (left side of the plot) are not stable across model orders. As higher-degree terms are added, previously estimated lower-order coefficients shift significantly, often changing sign and magnitude.
Figure 5(b) presents coefficient values for sparse regression using Legendre polynomials, with colors indicating contributions from different polynomial degrees. Here, a contrasting pattern emerges: coefficients associated with lower-degree terms remain stable as higher-degree terms are added. New coefficients primarily emerge in the higher-order region of the x-axis, without disturbing the existing ones. This stability results from the orthogonality of the Legendre basis, which decorrelates the polynomial terms and enables additive refinement without re-tuning existing coefficients.
(a) (a)
(b) (b)
Figure 6: (a) The scaling behavior of the sparse regression coefficients is the same as for coefficients of Fourier series. (b) The difference between the sparse regression fit and the real values for relative humidity of 5%.
The contrast between the Figs. 5(a) and (b) underscores the advantage of orthogonal polynomial bases in sparse regression. Simple regression results in unstable, entangled coefficient estimates that shift with basis expansion, complicating interpretability and reuse. Sparse regression with ordinary polynomial bases fails to converge for higher degrees. In contrast, sparsity and orthogonal polynomials yield stable, hierarchical models where lower-order structure is preserved and higher-order terms incrementally enrich the representation. This behavior is particularly valuable for symbolic regression and interpretable modeling, where each term ideally reflects a distinct, meaningful contribution to the model output.
In Fourier analysis, the magnitude of coefficients typically decays as $1/n$ (where $n$ is the order of the term) for functions of bounded variation [52] - a class that includes many naturally occurring signals and is a reasonable assumption for observational data. This decay reflects the fact that higher-order (or higher-frequency) components contribute less to the overall structure of such functions. A similar trend is observed in sparse regression using orthogonal polynomial bases, see Fig. 6(a). When coefficients are sorted by magnitude, they exhibit a clear decreasing pattern, analogous to the Fourier case, with lower-order terms capturing the dominant structure and higher-order terms refining the approximation in a controlled manner.
(a) (a)
(b) (b)
Figure 7: (a) Comparison of error histograms for the standard UTCI approximation and the sparse regression fit for a degree 10 Legendre polynomial. (b) Cumulative distribution for the absolute errors of the two models.
This suggests that through the use of sparse regression with an orthogonal polynomial basis we have achieved a Fourier-like decomposition of the UTCI Offset in the Legendre basis (instead of the trigonometric one). This has a number of theoretical advantages: due to the orthogonality of the basis functions, the decomposition minimizes the $L_{2}$ distance (least squares) between approximation and function, guaranteeing the best possible polynomial fit for a given model complexity [52]. Additionally, the coefficients are uncorrelated and hierarchically structured, ensuring that lower-order components remain stable as higher-order terms are added—enhancing both interpretability and numerical robustness. Fig. 6(b) shows the spatial distribution of the error between the proposed sparse regression fit (using a degree 10 Legendre polynomial) and the true UTCI Offset values for a fixed relative humidity of 5%. The errors are small and smoothly varying, indicating good agreement across the input space.
Fig. 7(a) presents a comparison of error histograms for both the standard approximation and the sparse model. The sparse model produces a narrower, more sharply peaked distribution centered at zero, highlighting a reduction in error variance and suggesting better generalization. Fig. 7(b) shows the cumulative distribution of absolute approximation errors for two models. The curve for our fit rises more steeply and reaches higher cumulative values at lower error thresholds, indicating that a larger proportion of predictions fall within smaller error margins. This reflects a clear improvement in accuracy: the sparse model not only reduces average error but also minimizes the frequency of large deviations compared to the standard approximation. For example, the frequency of absolute errors larger than $2^{\circ}$C is reduced from 8% to 4%, while the frequency of errors larger than $3^{\circ}$C reduced from 2% to 0.5%. These results confirm that the sparse regression approach can achieve comparable or improved predictive accuracy while maintaining interpretability and model parsimony.
4 Conclusion
This work has addressed key challenges in environmental and climate data modeling, with a focus on the Universal Thermal Climate Index (UTCI) as a complex, physiologically based measure of thermal comfort. We aimed at improving the approximation of the UTCI compared the to literature baseline. Accurately modeling such indices requires interpretable and efficient regression methods. Symbolic regression and grammar-based approaches—particularly when combined with sparse modeling—offer a powerful framework for discovering concise, physically meaningful relationships in data, enabling robust equation discovery in complex environmental systems.
Sparse regression methods, such as LASSO, helped reduce overfitting and improved interpretability. As we have shown, the choice of basis functions is crucial: orthogonal polynomials like Legendre polynomials offer better numerical stability and conditioning than monomials. They enable hierarchical models where higher-order terms don’t affect lower-order estimates, making them especially useful in sparse, interpretable models. Empirical results support these theoretical advantages. Using sparse regression with orthogonal polynomial basis we have:
  1. (a)
    Achieved better accuracy (lower loss with same number of parameters) than the state of the art in the literature.
  2. (b)
    Found a Pareto front for different model complexities. Loss curves reveal that sparse models with orthogonal bases efficiently populate a Pareto front, balancing complexity and accuracy.
  3. (c)
    Determined coefficients consistent over models with different capacities. Coefficient plots for models built on orthogonal bases show the progressive inclusion of higher-order components without disrupting lower-order structure, in contrast to models using simple regression and ordinary polynomials.
  4. (d)
    Trained the model over only 20% of the data with successful generalization in testing over the other 80%. The results are robust under bootstrapping.
  5. (e)
    Essentially decomposed the UTCI in a Fourier expansion with a Legendre-polynomial basis, with parameters scaling as expected. Thus, we are arguably close to the theoretical optimum results for a robust approximation in the $L_{2}$ metric (or least squares).
Additionally, error distributions for better models exhibit sharp, Gaussian-like peaks, reflecting low and consistent residuals. Visualizations of approximation error over a three-dimensional input space confirm this, with better models yielding tightly constrained error magnitudes across the domain. Together, these insights underscore the synergy between sparse modeling, orthogonal representations, and equation discovery methods in building interpretable, efficient models for complex environmental phenomena such as described by UTCI. Future work will extend this framework by integrating grammar-based equation discovery with sparse regression in orthogonal bases to uncover compact, physically consistent representations of different environmental indices and measurements.
The code to reproduce the results is available at:
https://doi.org/10.5281/zenodo.16880383
{credits}
4.0.1 Acknowledgements
This publication is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Postdoctoral Fellowship Programme, SMASH co-funded under the grant agreement No. 101081355. The operation (SMASH project) is co-funded by the Republic of Slovenia and the European Union from the European Regional Development Fund.
The authors acknowledge the financial support of the Slovenian Research Agency via the Gravity project AI for Science, GC-0001 and of the Slovenian Research And Innovation Agency (research core funding No. P1-0188).